OpenStudy (anonymous):

What is the solution set of (3x-1)^2=5?

OpenStudy (anonymous):

x = 1/3 (1-sqrt(5)) or x = 1/3 (1+sqrt(5)) Expanded Form: 9 x^2-6 x+1 = 5